OpenStudy (anonymous):

solve for x/3+5/2=10

OpenStudy (anonymous):

solve for \(x\)?

OpenStudy (anonymous):

if so, \[{x\over3}+{5\over2}=10\]get common denominators \((6)\):\[{2x\over6}+{15\over6}={60\over6}\]the denominators no longer matter:\[2x+15=60\]solve for x, subtract ](15\) from both sides:\[2x=45\]isolate the \(x\), divide both sides by \(2\):\[x={45\over2}\] or if you want it in decimal form:\[x=22.5\]

OpenStudy (anonymous):

x/3+5/√2=10 solve for x plz

OpenStudy (anonymous):

\[\large{{x\over3}+{5\over\sqrt{2}}}=10~??\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solve in simplest form not decimal plz

OpenStudy (anonymous):

follow the same steps as i gave above

OpenStudy (anonymous):

what do i do with the square root

OpenStudy (anonymous):

hint: the common denominator is \(3\sqrt{2}\)

OpenStudy (anonymous):

multiple by square root of 2?

OpenStudy (anonymous):

either way

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

quickly

OpenStudy (anonymous):

like first step .. what i should do

OpenStudy (anonymous):

\[{x\sqrt{2}\over3\sqrt{2}}+{15\over3\sqrt{2}}={30\sqrt{2}\over3\sqrt{2}}\]denominator no longer matters:\[x\sqrt{2}+15=30\sqrt{2}\]subtract 15:\[x\sqrt{2}=30\sqrt{2}-15\]\[x=\underline{~~~~~~~~~~~~~~~~~~~~}?\] hint:divide both sides by \(\sqrt{2}\)

OpenStudy (anonymous):

30-15 square of 2 over 2

OpenStudy (anonymous):

final answer ?